# Synthetic History: Evaluating Visual Representations of the Past in Diffusion Models

### v2

@@ 7 Discussion and Limitations
While our study offers the first systematic evaluation of historical representation in TTI models, it is constrained by several important limitations. First, our analysis of stylistic defaults relies on broad categories and does not capture finer distinctions. Second, while we selected a range of centuries and decades to capture broad trends and subtler shifts, these choices were somewhat arbitrary; alternative periods might yield different insights. Third, we based our prompts on universal activities, which could omit details relevant to specific historical contexts. Fourth, while our use of LLMs for demographic estimation and anachronism detection demonstrates promising initial results, we remain mindful of the challenges inherent in historical analysis, the potential biases highlighted in prior research (Gallegos et al., 2024; Navigli et al., 2023), and the fact that LLMs cannot replace the depth or reliability of expert historical knowledge. We are also aware that our baselines are significantly simplified, averaging demographic expectations across time and activity and omitting other factors. Similarly, the use of face classifiers further simplifies social complexities, reducing gender and race to discrete categories. Finally, while our anachronism pipeline was validated with a user study, judgments may vary due to differences in historical expertise and subjective interpretation. However, our approach establishes a strong basis for systematically evaluating historical aspects of TTI models, and its flexibility allows application to future models for continuous improvement.
@@ 8 Ethics statement
Our study includes human subjects for the purpose of evaluating anachronism detection. All annotations were carried out by adult participants recruited under fair working conditions and compensated at or above the local minimum wage. The task was limited to visual evaluation of synthetic images, and did not involve collection of personally identifiable information (PII) or sensitive data. Where required, the study received approval from the appropriate institutional ethics review process. Participants provided informed consent and were free to withdraw at any time.
@@
For demographic analysis, we employed an face classifier (FairFace) that treats race and gender as categorical attributes. We recognize that such classifications simplify complex, socially constructed identities and may reproduce existing biases. We therefore report demographic results in aggregate, and interpret them cautiously, emphasizing limitations and potential sources of misclassification. Our intention is not to essentialize or reify these categories, but to critically assess how generative models encode and reproduce demographic representation.
@@ 9 Reproducibility statement
We provide all scripts for reproducing the experiments, the trained linear style classifier, and a link to the generated images of the HistVis dataset with their metadata, available in an anonymous Kaggle repository as supplementary ZIP archives.
@@ Appendix I Anachronism Detection
Figure 13 plots the top 15 anachronistic elements per model, ranked by frequency and severity, and aggregated across different decades rather than centuries. Each point corresponds to an element proposed by the LLM and verified via VQA, with fuzzy matching used to unify lexical variations. Items like clothing continue to appear more frequently than they do severely (7-10%), whereas elements such as audio devices are comparatively rare but exhibit higher severity, once generated, they strongly conflict with the historical context. This pattern again illustrates that the models rely heavily on conceptual cues (e.g., “audio sources” for music-related prompts), often at the expense of accurate temporal conditioning.
@@
Figure 16 presents the user interface for the human evaluation study, which includes the task description given to participants and example questions. The task description introduces the concept of anachronisms to ensure participants understand it before beginning the task and provides illustrative examples of both historically accurate and anachronistic images for further clarification. Participants were then required to assess TTI-generated images for temporal inconsistencies by answering a series of yes-no questions.

### v3

@@ 7.1 Limitations
While our study offers the first systematic evaluation of historical representation in TTI models, it is important to emphasize several key limitations. First, our analysis of stylistic defaults is based on broad categories and therefore does not capture finer stylistic distinctions, which remains a potential direction for future work . Second, although we selected a range of different centuries, as well as decades, to capture broad trends and subtler shifts, these choices were somewhat arbitrary; alternative periods could have been chosen and might have yielded different insights. Third, in our benchmark we decided to focus on a broad spectrum of universal activities for this initial investigation, and we are aware that this may omit details relevant to specific historical contexts. Fourth, while our use of LLMs for demographic estimation and anachronism detection demonstrates promising initial results, we remain mindful of the challenges inherent in historical analysis, the potential biases highlighted in prior research (Gallegos et al., 2024; Navigli et al., 2023), and the fact that LLMs cannot replace the depth or reliability of expert historical knowledge. We are also aware that our baselines are significantly simplified, as they are obtained by averaging demographic expectations across period and activity and therefore omit other potentially relevant factors. Similarly, the use of face classifiers further oversimplifies social complexities by reducing gender and race to discrete categories. Finally, while our anachronism pipeline was validated with a user study, judgments may vary due to differences in historical expertise and subjective interpretation. Moreover, the LLM-generated anachronism list is not exhaustive; it targets prominent and recurring inconsistencies rather than every possible historical deviation. Despite these limitations, our approach establishes a strong inital basis for systematically evaluating historical aspects of TTI models, and its flexibility allows application to future models for continuous improvement.
7.2 Historical Plausibility and Diversity
Evaluating demographic patterns in historically-situated synthetic imagery raises an inherent ethical dilemma: how to generate content that accurately depicts the past while avoiding that such content further propagates biases and reinforces exclusionary narratives. Recent work on sociopolitical aspects of generated images (Shaw et al., 2025) suggests that users are more receptive to diversity when it is transparently and authentically contextualized, indicating that neither strict historical fidelity nor unrestricted diversification is sufficient on its own. Striking a balance between these competing aims remains an open challenge. Within this context, we find it necessary to emphasize that we position our LLM-derived historical plausibility as one possible evaluation axis, complementary to others such as cultural sensitivity, aesthetic quality, and faithfulness to the prompt, and not as an expected standard. Our intention is not to prescribe a “correct” way of depicting the past, but to provide a starting point for addressing this challenge at a broader, systematic scale.
8 Conclusion
We presented a benchmark for assessing how TTI diffusion models represent historical contexts. Our analysis is focused on three aspects: stylistic bias, chronological inconsistencies and demographic representations. Our results suggest that current models rely on narrow and stereotyped stylistic cues rather than nuanced understandings of historical periods, often defaulting to era-specific stylistic stereotypes and producing photorealistic depictions only from the 20th century onward. In addition, frequent anachronisms suggest that historical periods are not cleanly separated in the latent spaces of these models, since modern artifacts often emerge in pre-modern settings, undermining the reliability of TTI systems in education and cultural heritage contexts. Finally, our demographic analysis indicates that TTI outputs often diverge from LLM-estimated historically plausible expectations, sometimes aligning with contemporary expectations and sometimes reflecting activity-related biases rather than the specified historical condition. Ultimately, balancing historical fidelity with broader representational goals remains an open challenge, and future work should integrate richer historical knowledge and interdisciplinary expertise to develop models that depict the past with greater accuracy and sensitivity.
9 Acknowledgments
This research was supported by the Swiss National Science Foundation (SNSF) under the Ambizione Grant Scheme, Grant No. 216104.
@@ 10 Ethics statement
Our study includes human subjects for the purpose of evaluating anachronism detection. All annotations were carried out by adult participants recruited under fair working conditions and compensated at or above the local minimum wage. The task was limited to visual evaluation of synthetic images, and did not involve collection of personally identifiable information (PII) or sensitive data. Participants provided informed consent and were free to withdraw at any time.
@@
For demographic analysis, we employed a face classifier (FairFace) that treats race and gender as categorical attributes. We recognize that such classifications simplify complex, socially constructed identities and may reproduce existing biases. We therefore report demographic results in aggregate, and interpret them cautiously, emphasizing limitations and potential sources of misclassification. Our intention is not to essentialize or reify these categories, but to critically assess how generative models encode and reproduce demographic representation.
@@ 11 Reproducibility statement
Both the dataset111Hugging Face Repository, including all associated metadata, and our reproducible evaluation scripts222GitHub Repository are publicly available under cc by-nc 4.0 license.
@@ Appendix B HistVis Prompt Design
Prompt construction.
All prompts follow a predefined template (“a person [activity] in the [historical period]”), and the set of time periods was specified by the authors in advance. LLMs were used during the early ideation stage to surface candidate activities, which was then manually curated by the authors.
@@ Appendix I Anachronism Detection
Figure 13 plots the top 15 anachronistic elements per model, ranked by frequency and severity, and aggregated across different decades rather than centuries. Each point corresponds to an element proposed by the LLM and verified via VQA, with fuzzy matching used to unify lexical variations. Items like clothing continue to appear more frequently than they do severely (7-10%), whereas elements such as audio devices are comparatively rare but exhibit higher severity, once generated, they strongly conflict with the historical context. This pattern again illustrates that the models rely heavily on conceptual cues (e.g., “audio sources” for music-related prompts), often at the expense of accurate chronological conditioning.
@@
Figure 16 presents the user interface for the human evaluation study, which includes the task description given to participants and example questions. The task description introduces the concept of anachronisms to ensure participants understand it before beginning the task and provides illustrative examples of both historically accurate and anachronistic images for further clarification. Participants were then required to assess TTI-generated images for chronological inconsistencies by answering a series of yes-no questions.
@@
For this validation step, each LLM was prompted as follows:
For a given activity and historical period, estimate the plausible demographic breakdown based on global historical context:
Gender (%): Proportion of male and female participants.
Continent (%): Estimated distribution across the following regions: Africa, Asia, Europe, North America, South America, Oceania.
Base your estimates on historical population distributions, social structures, and participation norms relevant to the activity, and avoid assuming a single geographic focus.